OpenStudy (anonymous):

A rancher who wishes to fence off a rectangular area finds that the fencing in the east-west direction will require extra reinforcement due to strong prevailing winds. Fencing in the east-west direction will therefore cost $15 per (linear) yard, as opposed to a cost of $10 per yard for fencing in the north-south direction. The rancher wants to spend $7200 on fencing, express the area of the rectangle as a function of its width x. (by width we mean the measure in yards of a side running in the east-west direction.) Also, find Domain and find which width x yields the rectangle of largest area.

OpenStudy (anonymous):

width=x length = y

OpenStudy (anonymous):

so the total cost = 2 ( 15x + 10y ) =7200

OpenStudy (anonymous):

constraint ^

OpenStudy (anonymous):

then A= xy

OpenStudy (anonymous):

solve the constraint for y y= (7200 - 30x ) / 20

OpenStudy (anonymous):

sub into the area function

OpenStudy (anonymous):

A= (1/30 ) ( 7200x - 30x^2 )

OpenStudy (anonymous):

dA / dx = (1/30 ) (7200 -60x ) =0

OpenStudy (anonymous):

x=120

OpenStudy (anonymous):

d^2 A / dx^2 = -2 <0 , so the graph is concave down and this is a maximum

OpenStudy (anonymous):

whoops a few lines back it should be A= (1/20) ( 7200x -30x^2 ) , but that doesnt change the value of x where it is a maximum

OpenStudy (anonymous):

whoops a few lines back it should be A= (1/20) ( 7200x -30x^2 ) , but that doesnt change the value of x where it is a maximum

OpenStudy (anonymous):

and the second derivative is still negative